OpenStudy (samsan9):

find the vertices, foci, eccentricity, asymptotes, length of transverse axis, length of conjugate axis for (x-4)^2/36-(y-2)/9=1

OpenStudy (samsan9):

@cmorr23

OpenStudy (anonymous):

Alright find the center.

OpenStudy (samsan9):

(4,2)

OpenStudy (samsan9):

\[\frac{ (x-4)^2 }{ 36 }-\frac{ (y-2) }{ 9 }=1\]

OpenStudy (samsan9):

(y-2)^2

OpenStudy (anonymous):

Ok is it vertical or horizontal?

OpenStudy (anonymous):

Do you need help with that part?

OpenStudy (samsan9):

horizontal

OpenStudy (anonymous):

Yup good job. Ok so now where are your vertices?

OpenStudy (anonymous):

Remember they're going to be to the left and right of the center.

OpenStudy (anonymous):

So the y axis won't change.

OpenStudy (samsan9):

it is to the left and right but i dont by how much

OpenStudy (anonymous):

Ok so now to find the vertices you know how there are different numerators with the y in one and x in the other? You need to find the square root of the one with the x in the numerator since it is a horizontal equation. So find the square root of 36.

OpenStudy (samsan9):

6

OpenStudy (anonymous):

Good so the vertices are going to be 6 to the right and left of the center.

OpenStudy (anonymous):

So now what are your vertices?

OpenStudy (samsan9):

(10,2) and (-2,2)

OpenStudy (anonymous):

Yup good job. So now we'll find the foci which are represented by the equation of c^2=A^2-B^2. So which number is a^2 and which one is b squared?

OpenStudy (samsan9):

a^2= is 9 and B^2=36

OpenStudy (anonymous):

Oops my bad I meant c^2=A^2+B^2 because it's a hyperbola. But no it's the other way around Because A squared is always the bigger number and b squared is smaller. So A squared is equal to 36 and b squared is equal to 9. So now plug those into the equation and what do you get?

OpenStudy (samsan9):

45 and i think its square rooted

OpenStudy (anonymous):

Yeah so your foci are plus or minus the sqrt of 45.

OpenStudy (anonymous):

So do you know how to find eccentricity?

OpenStudy (samsan9):

isnt it c over a?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

So what's the eccentricity?

OpenStudy (anonymous):

sqaure root of 45 and then divide it by 6.

OpenStudy (samsan9):

i got 1.11803

OpenStudy (anonymous):

Yup and remember the eccentricity of a hyperbola is always greater than 1 so you can check your answer that way.

OpenStudy (anonymous):

Ok now for the the asymptotes what do you know about those?

OpenStudy (anonymous):

Lets do the traverse axis and conjugate axis they're really easy.

OpenStudy (samsan9):

ok show me how to find them

OpenStudy (anonymous):

Aight so for the traverse axis find the square root of A^2 and multiply it by 2 and you're done.

OpenStudy (anonymous):

And the same thing for the conjugate axis except you use b^2

OpenStudy (anonymous):

So what are your axes?

OpenStudy (samsan9):

so the transverse axis is 12 and the conjugate axis is 6

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Now for the asymptotes since your hyperbola isn't in the center you've gotta subract the y coordinates and x coordinates from b/a so it'll look like this. One sec.

OpenStudy (anonymous):

\[y +2=\pm3/4(x+4)\]

OpenStudy (anonymous):

\[y +2=\pm3/6(x+4)\]

OpenStudy (anonymous):

Once you work it out you get x-2y=0 and x+2y-8=0.

OpenStudy (samsan9):

so then what?

OpenStudy (anonymous):

Then you have to put them in slope intercept form and all that crap.

OpenStudy (samsan9):

wow so this is the long annoying part huh

OpenStudy (anonymous):

Yup I hate asymptotes. lol

OpenStudy (samsan9):

slope intercept form is the y=mx+b thing right?

OpenStudy (anonymous):

Yup THAT thing

OpenStudy (anonymous):

One sec i'll help with the asymptotes thing I just need to check over it again.

OpenStudy (samsan9):

ok

OpenStudy (anonymous):

I got it this time I wrote the equation wrong so the formula for asymptotes is y-k=b/a(x-h) With K=2 h=4 b=3 and a=6.

OpenStudy (anonymous):

So then you plug it in and it looks like this y-2=3/6(x-4).

OpenStudy (anonymous):

Next you can simplify 3/4 to being 1/2 then you multiple 1/2 times x-4 and you get y-2=1/2x-2

OpenStudy (anonymous):

Then you try to get the y by itself and add two to each side giving you y=1/2x because the 2's cancel each other out.

OpenStudy (anonymous):

Then to get your other asymptote you do the exact same thing except instead of having regular b/a you have negative b/a.

OpenStudy (anonymous):

Got it? Sorry I took so long i've been working all day.

OpenStudy (samsan9):

its cool thank you so much for all of the help

OpenStudy (anonymous):

Yup your welcome.

OpenStudy (samsan9):

do you know a website that could help me graph the two equations that i asked you to help me on?

OpenStudy (samsan9):

are you still there?

OpenStudy (anonymous):

Geogebra Is a good one that's what I used for my class it's not a website though it's a program.

OpenStudy (anonymous):

I think there's actually an online version of it too that you don't have to download. But I've only used the one you download it's free by the way.